# A uniform kernel trick for high and infinite-dimensional two-sample problems

### v2

@@ 6 Technical details
and hence $h_{\varepsilon}\to h^{+}$ in $\mathcal{H}_{k}$ (as $\varepsilon\to 0$). Now, we check that $\sigma_{k;\operatorname{P}-\operatorname{Q}}^{\prime}(g)=g\left(h^{+}\right)$, for $g\in\mathcal{C}\left(\mathcal{F}_{k},d_{\mathcal{H}_{k}}\right)$. We firstly observe that $h^{+}\in A_{\varepsilon}(\operatorname{P}-\operatorname{Q})$, for all $\varepsilon>0$. Hence, from equation (33), we have that $g\left(h^{+}\right)\leq\sigma_{k;\operatorname{P}-\operatorname{Q}}^{\prime}(g)$. On the other hand, we can extract a maximizing sequence $h_{m}\in A_{{}^{1}\!/_{m}}(\operatorname{P}-\operatorname{Q})$ ($m\in\mathbb{N}$) satisfying that $\underset{A_{{}^{1}\!/_{m}}(\operatorname{P}-\operatorname{Q})}{\operatorname{sup}}\,(g)\leq g\left(h_{m}\right)+\frac{1}{m}$. As $g$ is continuous and $h_{m}\to h^{+}$ as $m\to\infty$ in $\mathcal{H}_{k}$, we obtain that $\sigma_{k;\operatorname{P}-\operatorname{Q}}^{\prime}(g)=\underset{m\to\infty}{\operatorname{lim}}\,\underset{A_{{}^{1}\!/_{m}}(\operatorname{P}-\operatorname{Q})}{\operatorname{sup}}\,(g)\leq\underset{m\to\infty}{\operatorname{lim}}\,g\left(h_{m}\right)=g\left(h^{+}\right).$ Therefore, we obtain that $\sigma{k;\operatorname{P}-\operatorname{Q}}^{\prime}(g)=g\left(h^{+}\right)$, which is a linear mapping, so $\sigma_{k}$ is fully differentiable and the proof is complete. ∎
@@
where $\sigma_{k}$ is defined in (29). Using the same ideas as in the proof of (Cárcamo et al., 2020, Theorem 6.1), it can be checked that the paths of $\mathbb{G}$ in (37) are a.s. in $\mathcal{C}_{u}\left(\mathcal{F}_{k},\rho\right)$ (uniformly continuous), where
@@
is the natural $\operatorname{L}^{2}$-metric of $\mathbb{G}$. From (22), it can be readily checked that $\mathcal{C}_{u}\left(\mathcal{F}_{k},\rho\right)\subset\mathcal{C}_{u}\left(\mathcal{F}_{k},d_{\mathcal{H}_{k}}\right)$ and hence $\mathbb{G}\in\mathcal{C}\left(\mathcal{F}_{k},d_{\mathcal{H}_{k}}\right)$ a.s. To finish the proof it is enough to apply Lemma 4 together with the functional Delta method (van der Vaart and Wellner, 2023, Section 3.10). ∎
@@
Proof.
Let us fix $g\in\mathcal{C}\left(\mathcal{F}_{k,\Lambda},\rho\right)$. Again, from (Cárcamo et al., 2020, Theorem 2.1), we have that $\sigma_{k,\Lambda}$ is Hadamard directionally differentiable and
@@
Hence, from (44) and by taking $m\to\infty$ we obtain that $\lambda^{\star}\in\Lambda_{0}$. The fact that $d_{k,\Lambda}(\operatorname{P},\operatorname{Q})>0$ follows from 3 and the proof of ((i)) is complete.
To show ((ii)), using the same ideas as in the proof of equation (36) and ((i)), we obtain that
@@
Therefore, $\rho\left(h_{m},h^{+,\lambda_{m}}\right)\to 0$, and ((ii)) holds. To check ((iii)), by ((ii)), it is enough to see that $\rho\left(h^{+,\lambda_{m}},h^{+,\lambda^{\star}}\right)\to 0$, as $m\to\infty$. By (44) and repeatedly applying DCT (thanks to 3), it can be checked that
@@
Furthermore, for $m$ large enough, we have that
@@
To finish, we use (43), ((i)), ((iii)), as well as the continuity of the functional $g$ (with respect to the metric $\rho$) to obtain that
@@ Acknowledgments
This work has been partially supported by Grant PID2019-109387GB-I00 from the Spanish Ministry of Science and Innovation. The authors would like to thank two anonymous reviewers for their comments and references on the first version of this paper. We are also grateful to George Wynne (Imperial College, London) for useful comments regarding the dominance hypothesis for families of kernels in the infinite-dimensional case. We also want to thank Bojan Mihaljevic (Universidad Politécnica de Madrid) for his valuable assistance and advice in the development of the empirical section. Finally, kind help from Francisco de la Hoz (University of the Basque Country) with the final computations is also gratefully acknowledged. Some of the computational work has been done in the Center for Scientific Computation (UAM, project estfunc).

### v3

@@ 6 Technical details
and hence $h_{\varepsilon}\to h^{+}$ in $\mathcal{H}_{k}$ (as $\varepsilon\to 0$).
Now, we check that $\sigma_{k;\operatorname{P}-\operatorname{Q}}^{\prime}(g)=g\left(h^{+}\right)$, for $g\in\mathcal{C}\left(\mathcal{F}_{k},d_{\mathcal{H}_{k}}\right)$. We firstly observe that $h^{+}\in A_{\varepsilon}(\operatorname{P}-\operatorname{Q})$, for all $\varepsilon>0$. Hence, from equation (33), we have that $g\left(h^{+}\right)\leq\sigma_{k;\operatorname{P}-\operatorname{Q}}^{\prime}(g)$. On the other hand, we can extract a maximizing sequence $h_{m}\in A_{{}^{1}\!/_{m}}(\operatorname{P}-\operatorname{Q})$ ($m\in\mathbb{N}$) satisfying that $\underset{A_{{}^{1}\!/_{m}}(\operatorname{P}-\operatorname{Q})}{\operatorname{sup}}\,(g)\leq g\left(h_{m}\right)+\frac{1}{m}$. As $g$ is continuous and $h_{m}\to h^{+}$ as $m\to\infty$ in $\mathcal{H}_{k}$, we obtain that $\sigma_{k;\operatorname{P}-\operatorname{Q}}^{\prime}(g)=\underset{m\to\infty}{\operatorname{lim}}\,\underset{A_{{}^{1}\!/_{m}}(\operatorname{P}-\operatorname{Q})}{\operatorname{sup}}\,(g)\leq\underset{m\to\infty}{\operatorname{lim}}\,g\left(h_{m}\right)=g\left(h^{+}\right).$ Therefore, we obtain that $\sigma{k;\operatorname{P}-\operatorname{Q}}^{\prime}(g)=g\left(h^{+}\right)$, which is a linear mapping, so $\sigma_{k}$ is fully differentiable and the proof is complete. ∎
@@
where $\sigma_{k}$ is defined in (29).
Using the same ideas as in the proof of [14, Theorem 6.1], it can be checked that the paths of $\mathbb{G}$ in (37) are a.s. in $\mathcal{C}_{u}\left(\mathcal{F}_{k},\rho\right)$ (uniformly continuous), where
@@
is the natural $\operatorname{L}^{2}$-metric of $\mathbb{G}$. From (22), it can be readily checked that $\mathcal{C}_{u}\left(\mathcal{F}_{k},\rho\right)\subset\mathcal{C}_{u}\left(\mathcal{F}_{k},d_{\mathcal{H}_{k}}\right)$ and hence $\mathbb{G}\in\mathcal{C}\left(\mathcal{F}_{k},d_{\mathcal{H}_{k}}\right)$ a.s. To finish the proof it is enough to apply Lemma 4 together with the functional Delta method [48, Section 3.10]. ∎
@@
Proof:.
Let us fix $g\in\mathcal{C}\left(\mathcal{F}_{k,\Lambda},\rho\right)$. Again, from [14, Theorem 2.1], we have that $\sigma_{k,\Lambda}$ is Hadamard directionally differentiable and
@@
Hence, from (44) and by taking $m\to\infty$ we obtain that $\lambda^{\star}\in\Lambda_{0}$. The fact that $d_{k,\Lambda}(\operatorname{P},\operatorname{Q})>0$ follows from 3 and the proof of 6 is complete.
To show 6, using the same ideas as in the proof of equation (36) and 6, we obtain that
@@
Therefore, $\rho\left(h_{m},h^{+,\lambda_{m}}\right)\to 0$, and 6 holds.
To check 6, by 6, it is enough to see that $\rho\left(h^{+,\lambda_{m}},h^{+,\lambda^{\star}}\right)\to 0$, as $m\to\infty$. By (44) and repeatedly applying DCT (thanks to 3), it can be checked that
@@
Furthermore, for $m$ large enough, we have that for $\operatorname{S}\in\{\operatorname{P},\operatorname{Q}\}$
@@
To finish, we use (43), 6, 6, as well as the continuity of the functional $g$ (with respect to the metric $\rho$) to obtain that
@@ Acknowledgments
This work has been partially supported by Grant PID2019-109387GB-I00 from the Spanish Ministry of Science and Innovation. A. Cuevas acknowledges as well financial support from Grant CEX2019-000904-S funded by MCIN/AEI/10.13039/501100011033. L.A. Rodríguez acknowledges funding from the German Science Foundation DFG Research Unit 5381: “Mathematical Statistics in the Information Age–Statistical Efficiency and Computational Tractability”. The authors would like to thank two anonymous reviewers for their comments and references on the first version of this paper. We are also grateful to George Wynne (Imperial College, London) for useful comments regarding the dominance hypothesis for families of kernels in the infinite-dimensional case. We also want to thank Bojan Mihaljevic (Universidad Politécnica de Madrid) for his valuable assistance and advice in the development of the empirical section. Finally, kind help from Francisco de la Hoz (University of the Basque Country) with the final computations is also gratefully acknowledged. Some of the computational work has been done in the Center for Scientific Computation (UAM, project estfunc).